OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = 3/sqrt3x+4

OpenStudy (anonymous):

\[y=\frac{ 3 }{ \sqrt{3x+4} }\]

OpenStudy (anonymous):

@jim_thompson5910 can you help? Please:)

jimthompson5910 (jim_thompson5910):

are you given answer choices?

jimthompson5910 (jim_thompson5910):

because this question is open-ended (ie there are multiple ways to do this)

OpenStudy (jdoe0001):

off the multiple choices as jim_thompson5910 already pointed out think about if you were to make f(x) = 3/x what will be your g(x)?

OpenStudy (anonymous):

f(x) = \[\frac{ 3 }{ \sqrt{x} }\], g(x) = 3x + 4 f(x) =\[\sqrt{3x+4}\] , g(x) = 3 f(x) =\[\frac{ 3 }{ x }\] , g(x) = 3x + 4 f(x) = 3, g(x) = \[\sqrt{3x+4}\]

OpenStudy (anonymous):

Sorry thats really hard to read but those are the answer choices

OpenStudy (anonymous):

A) f(x) =3/sqrt x , g(x) = 3x + 4 B) f(x) = sqrt 3x+4 , g(x) = 3 C) f(x) = 3/x , g(x) = 3x + 4 D) f(x) = 3, g(x) = sqrt 3x+4

OpenStudy (anonymous):

Answer choices!

jimthompson5910 (jim_thompson5910):

if you look at choice C, saying f(g(x)) would get you this f(x) = 3/x f(g(x)) = 3/(g(x)) ... replace every x with g(x) f(g(x)) = 3/(3x+4) ... replace every g(x) on the right with 3x+4 so this is close to what we want, but not quite

jimthompson5910 (jim_thompson5910):

so C is out

OpenStudy (anonymous):

Okay!

jimthompson5910 (jim_thompson5910):

see how I got that?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so what is f(g(x)) for choice B?

OpenStudy (anonymous):

\[\sqrt{13}\]

jimthompson5910 (jim_thompson5910):

so that's out too because that's not what we want

jimthompson5910 (jim_thompson5910):

for choice D, f(x) is always 3, so no matter what you plug in for x, it's going to remain 3 besides, there is no x to replace anyway (on the right side) so for choice D, f(g(x)) = 3 so D is out

OpenStudy (anonymous):

Okay, so the answer is a

jimthompson5910 (jim_thompson5910):

yep, I would do the same thing we've been doing to check though (so you get the practice)

OpenStudy (anonymous):

okay, Thank you so much!

jimthompson5910 (jim_thompson5910):

yw